OpenStudy (anonymous):

x^2 - y^2 = 7, x =4 <-- region is bounded by these two curves. Rotate this region about the y = 5 and find the volume.

OpenStudy (anonymous):

Using shell method

OpenStudy (anonymous):

MY ATTEMPT: r = (4 - y) h = (3 - \sqrt{y^2 + 7})

OpenStudy (anonymous):

you love these volume problems eh?

OpenStudy (anonymous):

I know this is hyperbola

OpenStudy (anonymous):

With the x-int's being (+/- root{7},0)

OpenStudy (anonymous):